OpenStudy (anonymous):

Which of the expressions are NOT polynomials? Check all that apply. A. x^2-(square root of x)-3 B. 4 - 3x + 5x^6 C. 5x^1/2 + 4x^2 D. 14x^7 E. x^2-x-12/x+3 F. 8x^10 + 2x^5

OpenStudy (saifoo.khan):

i think D.

OpenStudy (saifoo.khan):

and others.

jimthompson5910 (jim_thompson5910):

D is a polynomial, all monomials are polynomials

OpenStudy (saifoo.khan):

isn't D a monomial?

OpenStudy (saifoo.khan):

why?

jimthompson5910 (jim_thompson5910):

a monomial is a polynomial

OpenStudy (saifoo.khan):

O.O

jimthompson5910 (jim_thompson5910):

a polynomial is the sum of one OR more monomials

jimthompson5910 (jim_thompson5910):

it's weird, but you can technically sum one item

OpenStudy (saifoo.khan):

so, monomialS are polynomial.

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (saifoo.khan):

xD

jimthompson5910 (jim_thompson5910):

to answer the original question, the answers are A, C, and E

OpenStudy (saifoo.khan):

why are the answers A, C & E ?????????

jimthompson5910 (jim_thompson5910):

they want to know which ones are NOT polynomials

OpenStudy (saifoo.khan):

NOT a polynomial means a monomial right? so why are they monomial?

jimthompson5910 (jim_thompson5910):

NOT a polynomial would be something like \[\large \sqrt{x}\] or \[\large \frac{1}{x}\]

jimthompson5910 (jim_thompson5910):

again, monomials are polynomials

OpenStudy (saifoo.khan):

1/x can't be a monomial.... :O as you said. it can be written as x^-1. soo?

jimthompson5910 (jim_thompson5910):

exactly, 1/x is not a monomial and 1/x is not a polynomial

OpenStudy (saifoo.khan):

wow, then what is it?

jimthompson5910 (jim_thompson5910):

1/x is a rational expression, something completely different

jimthompson5910 (jim_thompson5910):

not sure where the confusion lies, but a polynomial is a very strict form that was explained in that other post

OpenStudy (saifoo.khan):

a^2 + a is that a polynomial?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (saifoo.khan):

E. x^2-x-12/x+3 why is that NOT a polynomial?

jimthompson5910 (jim_thompson5910):

didn't see that one correctly and forgot that you can simplify So because (x^2-x-12)/(x+3) = x-4, this means that it is a polynomial However, something like (x^2-x-12)/(x+9) cannot be simplified and isn't a polynomial

OpenStudy (saifoo.khan):

Alright, thanks. i think i got it. :D